OpenStudy (anonymous):

How can I do #20 without going through each option and doing out the series?

OpenStudy (anonymous):

Attached is a picture of the problem

OpenStudy (whpalmer4):

Well, you'd need to be familiar with the series expansion for the building blocks you've got there in those expressions. Do you know the expansion for e^x or e^(x^2)? How about sin x and cos x? Do any of them look similar in form to the mystery function?

OpenStudy (anonymous):

I know that e^x is 1 + x/2 + x^2/6 ...etc. right? is that what you mean by building blocks?

OpenStudy (whpalmer4):

Indeed! Now what if you multiplied that series by x^2? Doesn't it look like the mystery function, except with a pesky x^2 and x^3 term at the front? By the way, isn't it 1 + x + x^2/2 + x^3/6 + x^4/24 ...?

OpenStudy (anonymous):

yeah it is, I forgot the x. hmm i have a feeling it is choice D

OpenStudy (whpalmer4):

That's an excellent feeling :-) Multiplying the sequence gets you x^2 + x^3 + x^4/2 + x^5/6 (= x^5/3!) + x^6/24 (=x^6/4!) etc....then subtract x^2 and x^3 and there's your mystery function.

OpenStudy (anonymous):

perfect thanks :). could you remind me what the building blocks are for sin(x) and cos(x)?